OpenStudy (anonymous):

The length of a rectangle is 7 inches more than the width. The perimeter is 42 inches. Find the length and width

OpenStudy (anonymous):



OpenStudy (anonymous):

Cool, so for that rectangle, perimeter is 42. How do you find the perimeter?

OpenStudy (anonymous):

Do you know the formula for finding the perimeter of an object?

OpenStudy (anonymous):

i think its p=2l+2w

OpenStudy (anonymous):

yeah, in this case for the rectangle, P = 2*length + 2*width. So you have values for the length and the width, right? L = x + 7 W = x

OpenStudy (anonymous):

Plug those values into the perimeter function, what does the equation look like after doing that?

OpenStudy (anonymous):

my equation looks like this 42=2(7)+2w 42-14=28 which makes w = 14 then for L 14+7=21 am i right

OpenStudy (anonymous):

Getting there, so the values you have for each dimension are actually variables - (x+7), and x Therefore P = 2*L + 2*W becomes L = x + 7 W = x P = 2*(x+7) + 2*(x) 42 = 2*(x+7) + 2*(x)

OpenStudy (anonymous):

Because there is only one variable, one unknown, that being 'x', we can easily then solve for it from the adjusted equation.

OpenStudy (anonymous):

Does that make sense? Give that equation a shot, and solve for x.

OpenStudy (anonymous):

x=7

OpenStudy (anonymous):

Definitely! that's it, so L = ? and W = ?

OpenStudy (anonymous):

Just have to do that last step and you're donw

OpenStudy (anonymous):

Quite easy after finding x...

OpenStudy (anonymous):

so w = 7 and l=14

OpenStudy (anonymous):

Nice work

OpenStudy (anonymous):

thanks